OpenStudy (anonymous):

A population of scores has σ=10. in this population, a score of X=20 corresponds to z=-0.40. what is the population mean?

OpenStudy (anonymous):

I think z=(mean-X)/ò. Sorry I can't be bothered to find that symbol So -0,4=(mean-20)/10 So mean = 16 ?